OpenStudy (here_to_help15):

@mathmath333 @jim_thompson5910 @nincompoop

OpenStudy (haichi):

Sorry @Here_to_Help15 Not at that stage of math yet :(

OpenStudy (here_to_help15):

Oh its ok thanks

OpenStudy (here_to_help15):

@mathmath333 @jim_thompson5910

OpenStudy (here_to_help15):

sorry if i am overwhelming you :(

OpenStudy (here_to_help15):

i got AC=32

OpenStudy (mathmath333):

how do u got that

OpenStudy (here_to_help15):

did you get that* and hold on

OpenStudy (here_to_help15):

DF is to DE as AC is EC or DF/DE=AC/EC or numerically, 16/19=x/38 Where DF=16, DE=19 which is 1/2 of EC (midpoint), X is labeled for AC, and EC is given as 38. solve for x=AC=608/19=32. This i what i put down but i need to make it brief and organized thats what i need help on

OpenStudy (mathmath333):

lol we did this type of problem earlier

OpenStudy (mathmath333):

yes its \(\Huge \tt \begin{align} \color{black}{\checkmark }\end{align}\)

OpenStudy (here_to_help15):

Yes i know i am correct lol i need your help on making it brief and organized

OpenStudy (here_to_help15):

How would i do this

OpenStudy (here_to_help15):

@mathmath333 ?

OpenStudy (mathmath333):

u need to show that \(\large\tt \begin{align} \color{black}{\triangle EDF \approx \triangle ECA }\end{align}\) as \(ED \approx EC\) \(EF \approx EA\) by SS theorm

OpenStudy (mathmath333):

or simpliy bu this DF/DE=AC/EC

OpenStudy (here_to_help15):

bu?

OpenStudy (mathmath333):

by

OpenStudy (here_to_help15):

At @mathmath333 i need a explanation from you if you don't it seems simpler if that's alright with you?

OpenStudy (here_to_help15):

mind*

OpenStudy (mathmath333):

do u agree DF/DE=AC/EC

OpenStudy (here_to_help15):

Yes i agree with that

OpenStudy (mathmath333):

so by that we can say that trinagle EDF is similar to triangle ECA

OpenStudy (mathmath333):

do u still agree

OpenStudy (here_to_help15):

yes since they are similar correct?

OpenStudy (here_to_help15):

In other words i agree @mathmath333 i see what you are saying

OpenStudy (mathmath333):

wait u need to show first that \(\dfrac{ED}{EC}=\dfrac{EF }{EA}\) CAN U SHOW THAT

OpenStudy (here_to_help15):

Whats with the caps lol and what do you mean by can i show that

OpenStudy (here_to_help15):

Yes i can type that out if that's what you are saying

OpenStudy (mathmath333):

can u prove that

OpenStudy (here_to_help15):

Hm let me see

OpenStudy (here_to_help15):

Oh c is g i believe

OpenStudy (here_to_help15):

If we look at the question it doesn't show C

OpenStudy (mathmath333):

show C?

OpenStudy (here_to_help15):

No lol C is suppose to be G correct? since C is not in the problem

OpenStudy (mathmath333):

where is G?

OpenStudy (here_to_help15):

OM lol G is SUppose to be C

OpenStudy (here_to_help15):

I mean C IS SUPPOSE TO BE G

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{ED}{EC}\\ =\dfrac{ED}{ED+DC}\\ =\dfrac{ED}{ED+ED}\\ =\dfrac{ED}{ED+ED}\\ =\dfrac{1}{2}\\ }\end{align}\)

OpenStudy (mathmath333):

similarly u have to show that \(\large\tt \begin{align} \color{black}{\dfrac{EF}{EA}\\ =\dfrac{1}{2}\\ }\end{align}\)

OpenStudy (here_to_help15):

Look at the problem there is no C

OpenStudy (here_to_help15):

C is G correct?

OpenStudy (mathmath333):

it is C not G

OpenStudy (here_to_help15):

Where does it show C? in the problem

OpenStudy (mathmath333):



OpenStudy (mathmath333):

it is called C

OpenStudy (here_to_help15):

Ok nvm :P continue

OpenStudy (here_to_help15):

I just need to sum what i said up

OpenStudy (mathmath333):

yea nm it happens

OpenStudy (here_to_help15):

But briefly and organize

OpenStudy (here_to_help15):

I want to eat then hit the bed lol

OpenStudy (mathmath333):

so we proved that \(\dfrac{ED}{EC}=\dfrac{EF }{EA}\)

OpenStudy (here_to_help15):

Yes

OpenStudy (mathmath333):

hence \(\large\tt \begin{align} \color{black}{\triangle EDF \approx \triangle ECA }\end{align}\)

OpenStudy (here_to_help15):

yes

OpenStudy (mathmath333):

so \(\dfrac{ED}{EC}=\dfrac{DF }{CA}\)

OpenStudy (here_to_help15):

Ok continue

OpenStudy (mathmath333):

lol

OpenStudy (here_to_help15):

Oh this irrelevant but i will beat ya at chess

OpenStudy (here_to_help15):

Anyways continue lol

OpenStudy (mathmath333):

really

OpenStudy (here_to_help15):

yes really lol

OpenStudy (mathmath333):

ok thats it u need to show

OpenStudy (here_to_help15):

Ok but how @mathmath333 put it this way how would you answer this briefly and organized fit it all in 1 post please

OpenStudy (mathmath333):

i did it in organized way

OpenStudy (here_to_help15):

Lol fit it in 1 post please

OpenStudy (here_to_help15):

put everything together

OpenStudy (mathmath333):

after u i showed my last post continue from here DF is to DE as AC is EC or DF/DE=AC/EC or numerically, 16/19=x/38 Where DF=16, DE=19 which is 1/2 of EC (midpoint), X is labeled for AC, and EC is given as 38. solve for x=AC=608/19=32. This i what i put down but i need to make it brief and organized thats what i need help on

OpenStudy (here_to_help15):

ugh :/

OpenStudy (here_to_help15):

Im just going to turn it in my head hurts :/